OpenStudy (anonymous):

Derive the equation of the parabola with a focus at (4, −7) and a directrix of y = −15. Put the equation in standard form. Answer choices: a) f(x) = 1/16 x^2 -8x +11 b) f(x) = 1/16 x^2 -8x -10 c) f(x) = 1/16 x^2 -1/2x +11 d) f(x) = 1/16 x^2 -1/2x -10

OpenStudy (anonymous):

11?

OpenStudy (anonymous):

I mean 4,11?

OpenStudy (michele_laino):

Please, if I apply the definition of a parabola, namely "a parabola is a curve for which the distance between its point and focus is equal to the distance between that same point and the directrix", I can write: \[(x-4)^{2}+(y+7)^{2}=(y+15)^{2}\] Now, please continue the computation above

OpenStudy (michele_laino):

Please note that I called (x,y) a generic point which belongs to our parabola

OpenStudy (anonymous):

lolwhat how do u solve that

OpenStudy (michele_laino):

for example: \[(x-4)^{2}=x ^{2}+16-8x\] please do the same with the remaining terms

OpenStudy (anonymous):

ah ok so we're foiling them ?

OpenStudy (michele_laino):

we have to develop the squares

OpenStudy (anonymous):

(y+7)^2= y^2 + 14y + 49 ?

OpenStudy (anonymous):

(y+15)^2 = y^2 + 30y + 225?

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

yay :)

OpenStudy (anonymous):

what do I do from here?

OpenStudy (michele_laino):

I insert your result into my first equation, namely: \[x ^{2}+16-8x+y ^{2}+49+14y=y ^{2}+225+30y\] Now, please simplify

OpenStudy (anonymous):

are we solving for x or y

OpenStudy (michele_laino):

we have to solve for y

OpenStudy (anonymous):

does y= x^2 over 16 - x/2 -10 ?

OpenStudy (michele_laino):

perfect! Well done! :)

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

how do I find my answer with this?

OpenStudy (michele_laino):

please note that your "y" stands for "f(x)", so?

OpenStudy (anonymous):

yeah but there is no x^2 over 16 in the answer choices :(?

OpenStudy (michele_laino):

Please, note that your option is d), namely the fourth option

OpenStudy (michele_laino):

since 1/16 x^2 is the same as x^2 over 16

OpenStudy (anonymous):

Oh I see now thank you very much

OpenStudy (michele_laino):

thank you!